OpenStudy (anonymous):

Find all solutions of the equation algebraically: csc x + cot x = 1

OpenStudy (anonymous):

30°, 150° and 270°

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

csc(x) + cot(x) = 1 1/sin(x) + cos(x)/sin(x) = 1 1 + cos(x) = sin(x) sin(x) - cos(x) = 1 Square both sides sin²(x) - 2sin(x)cos(x) + cos²(x) = 1 1 - 2sin(x)cos(x) = 1 sin(2x) = 0 2x = {0, π, 2π, 3π} x = {0, π/2, π, 3π/2} Check csc(0) + cot(0) is undefined csc(π/2) + cot(π/2) = 1 csc(π) + cot(π) is undefined csc(3π/2) + cot(3π/2) = -1 The only value of x that satisfies this equation in the interval [0, 2π) is x=π/2

OpenStudy (anonymous):

primary lesson Trigono... :)

OpenStudy (anonymous):